OpenStudy (anonymous):

how do graph 2x-3y=12

OpenStudy (anonymous):

First you need to put it into slope-intercept form which is y=mx+b m is the slope and b is the y intercept

OpenStudy (anonymous):

2x-3y=12 -3y=12-2x y=-4+(2/3)x

OpenStudy (anonymous):

Your slope is -4/1 and your y intercept is 2/3

OpenStudy (anonymous):

To graph it start by finding the y intercept. once you have the point of (0,2/3) then you can use the slope to find the rest of your points

OpenStudy (anonymous):

Make sense

OpenStudy (anonymous):

since you have y=2/3x+4 should you start at 4 on the y axis and then do the slope over 2 up 3

OpenStudy (anonymous):

woops sorry I had it back wards yours is correct so your slope is actually 2/3 and your y intercept is (0,-4)

OpenStudy (anonymous):

i thought they y intecept was 4?

OpenStudy (anonymous):

It's -4 wrote it wrong sorry your problem should look like this now: y=2/3x-4

OpenStudy (anonymous):

So y intercept is (0,-4) and to find your other points remember that slope is rise, over run just like how an airplane has to get the speed before it can fly. so you will go over 2 and up 3.

OpenStudy (anonymous):

i see what i was doing wrong, yours was right. thanks

OpenStudy (anonymous):

Sorry about that... No problem